OpenStudy (anonymous):

Laura insured her car for $12,000. When it was damaged she got $8,000 from the insurance company. What percentage of the insured amount did Laura get?

OpenStudy (dumbcow):

8/12 = 2/3 = .6667 --> 66.67%